# Reflection-Reinforced Self-Training for Language Agents

### v1

@@ 3.2 Sequential Decision-Making
We also assess the proposed approach on sequential decision-making using ALFWorld (Shridhar et al., 2020). ALFWorld comprises a collection of text-based settings designed to test an agent’s ability to complete multi-step tasks across diverse interactive environments. Following Yao et al. (2022); Shinn et al. (2023), we operate under the assumption that the agents are devoid of any access to successful trajectories, relying solely on a binary indicator of task success or failure. Our evaluation encompasses testing the agent across 134 previously unseen environments, spanning six diverse tasks. These tasks range from locating concealed items and transporting objects to interacting with objects using other items.
| Model | Train Data Used (%) | Sucess Rate (%) |
| Llama-7B | - | 8.9 |
| Llama-7B w/ Self-Training | 11.2 | 37.3 |
| Llama-7B w/ Re-ReST | 48.0 | 51.4 |
@@
We build both the agent model and the reflector upon the Llama2-7b (Touvron et al., 2023). At each step, the agent can either contemplate its next move or generate admissible actions for execution as in Yao et al. (2022). Following the heuristics outlined by Shinn et al. (2023), we trigger the reflector model for self-reflection if the agent repeats an action with the same response over three cycles, or if it performs over 30 actions in an environment.
@@
We use one-shot prompting instead of the two-shot prompting in Shinn et al. (2023) for the models so that we can better fit a trajectory into the context window of Llama-2. We train the agent and reflector models on the collected trajectories for 2 epochs with a learning rate of 2e-5 using LoRA. Please refer to the appendix for the prompt template utilized during the inference process.
@@ Results.
We present the results in Table 2. It is evident that the base Llama model faces challenges in adapting to the experimental environment, but self-training can significantly improve the model performance. A significant point to highlight is that the model operates without access to complete trajectories during the experiment. Despite this limitation, it demonstrates a notable improvement in performance within unseen environments—increasing the success rate from 8.9% to 37.3% through the utilization of self-augmented trajectories. Furthermore, the implementation of the reflector contributes an 14.1% uplift in success rates, which affirms the efficacy of our proposed method.
@@ 3.3 Programming: Code Generation and Visual Question Answering
For code generation, we experiment with the Python code writing task on MBPP (Austin et al., 2021) and visual programming on GQA (Hudson & Manning, 2019). The MBPP benchmark consists of around 1,000 Python programming problems, with each problem paired with unit test cases. We follow its official split for the training and test data. The availability of the training set and its provided unit test cases make it suitable for our reflector to reflect and correct the model-generated code. For GQA, we randomly sample a subset of 5,000 data points for training and 1,000 data for testing.
@@
We build both the agent model and the reflector upon the CodeLlama-13B model (Roziere et al., 2023), which is the Llama-2 model finetuned for code generation. For MBPP, following Roziere et al. (2023), the agent model is given the unit test cases during code generation. Similarly, the reflection model is given the agent generation and its unit test results as the environmental feedback, and then generates a corrected version. For GQA, following Surís et al. (2023), we build the agent by providing a pre-defined set of visual APIs (e.g. object detection) and prompt the model to generate code to answer the question using the APIs.
@@
For MBPP, we use zero-shot and three-shot prompting for zero-shot agent and reflector data generation respectively. For GQA, we follow the prompt in Surís et al. (2023) for the model for sample generation. For both datasets, the agent model samples 3 generations per training instance as before. We do not use the provided ground truths for MBPP training for consistency with the other experimental settings. The agent and reflector models are trained for 3 epochs with a learning rate of 3e-4 using LoRA.
@@
| Train Data Used (%) | Pass @1 (%) | Train Data Used (%) | Score (%) | |
| CodeLlama-13B | - | 48.6 | - | 40.9 |
| CodeLlama-13B w/ Self-Training | 66.9 | 54.5 | 44.7 | 41.9 |
| CodeLlama-13B w/ Re-ReST | 77.3 | 56.4 | 55.7 | 42.6 |
@@ Results.
As shown in Table 3, for MBPP, because CodeLlama is trained on a large amount of code generation corpus, the base CodeLlama model can achieve a decent performance without any finetuning. The high pass rate results in many of the training instances being used for self-training. After self-training on the MBPP training data, the model performance can be improved from 48.6% to 54.5%. The reflector model can generate more self-training data and the pass rate can be improved with the reflector-generated data. For GQA, similar improvements can be seen, indicating that our method is also applicable in the visual programming domain.
@@ 3.4 Text-to-Image Generation
| Model | Train Data Used (%) | VPEval Skill Score (%) | ||
@@
| Model | Train Data Used (%) | Dev. Set EM (%) |
@@
| Self-Consistency (6 agent samples) | 30.8 |
| Self-Consistency (3 agent samples + 3 reflector samples) | 32.0 |
| Model | HotpotQA EM (%) | MBPP Pass@1 (%) | VPEval Score (%) | ||
@@
We use the VPGen model proposed in Cho et al. (2023) as our base model, which is based on Vicuna-13B (Chiang et al., 2023) and is finetuned for text-to-layout generation on multiple constructed image-text datasets. The generated layouts are fed into an external model (i.e., GLIGEN (Li et al., 2023b)) for image generation. We build both the agent and reflector upon the VPGen model.
@@
We use the VPGen model to perform inference on their training data, and evaluate the generations using the VPEval framework (Cho et al., 2023). Specifically, during evaluation, a visual question answering model (BLIP-2 (Li et al., 2023a)) is used to determine if the generated images correctly capture the input text information. The BLIP-2 generated results are treated as the environmental feedback for the reflector. We do not use zero-shot reflection results to train the reflector because LLMs cannot perform this task. The agent and reflector models are trained for 2 epochs with a learning rate of 1e-5 using LoRA.
@@ 3.5 Analysis
In this part, we perform analyses to provide insights into our designs. The experiments are on HotpotQA unless otherwise noted.
@@
In this part, we investigate if we can simply sample more generations from the language agent for self-training and achieve comparable performance with our reflector-augmented method. Specifically, we try to sample generations for each instance, where is set to , and use the generated samples for self-training. As shown in Figure 2, if we keep sampling more generations from the language agent, the agent can indeed solve more instances and we can obtain an increasing amount of data for self-training. However, 1) the number of solved instances is still lower than the number of reflector-solved instances, demonstrating that the reflector can find the correct solutions more efficiently than sampling; 2) the model performance is not always improved with more training data and it cannot outperform our method even when trained with more generated samples, indicating that the quality of the self-training data is also important and our reflector can generate training data effectively for the agent.
@@
As illustrated in the method section, we propose to first train the reflector before using it to generate the self-training data. In this part, we investigate if we can use the reflector to perform self-correction in a zero-shot manner and then train the language agent. As in Table 5, we find that while the reflector can perform self-correction without any finetuning and improve the performance of the language agent, further improvements can be made if we specifically train the model for self-correction, demonstrating the effectiveness of our proposed reflector training strategy.
Re-ReST with Self-Consistency.
Previously, our reflector only functions during training and is not used during inference because it is often impossible to tell if a predicted answer is correct. Self-consistency (Wang et al., 2023a) is a decoding technique that combines multiple model predictions. It first samples multiple reasoning paths and then selects the most consistent answer by taking a majority vote, enabling us to also apply the reflector during inference. Specifically, we can first sample multiple answers from our model and then perform reflection with each output, regardless of whether the answer is correct. Then, we aggregate all the answers with self-consistency. As shown in Table 6, our method (3 agent samples and 3 reflection samples) can be integrated with self-consistency and achieve improvements over baseline (self-consistency with 6 model samples), indicating the potential of our method during inference.
@@
Our reflector turns incorrect samples into correct ones, naturally making negative-positive pairs that are suitable for preference optimization objectives such as DPO. In this part, we investigate the application of DPO in our method. As in Table 7, integrating DPO into our method can generally improve or achieve comparable performance with training models only with supervised training on positive samples, indicating the compatibility of Re-ReST with DPO.
@@ 4 Related Work
Recently, there has been a significant surge in research focusing on self-improving algorithms for language modeling. In this section, we first overview the research progress in language agents, then briefly describe self-training and self-correction methods for improving language agents.
@@
Language agents refer to language models that interact with the world in general. It has been demonstrated that LLMs can perform actions by generating specific commands (Nakano et al., 2021; Huang et al., 2022; Ahn et al., 2022) and calling external tool APIs (Lu et al., 2024; Schick et al., 2024; Gou et al., 2024). By integrating the model reasoning and acting abilities, ReAct (Yao et al., 2022) asks an LLM to first generate reasoning traces and then act accordingly, which is then improved by follow-up works through inference-time techniques such as reflection (Shinn et al., 2023) and planning (Yao et al., 2023). Recently, finetuning agents (Chen et al., 2023; Yin et al., 2024) have attracted attention from the research community. However, most of the existing works attempt to distill knowledge from a relatively strong LLM (e.g., GPT-4) to a weaker LLM (e.g., LLaMa-2). By contrast, our work bootstraps a language agent’s performance by utilizing its own reflective ability without using external models.
@@
Various self-training algorithms have been proposed to improve language models (He et al., 2019; Huang et al., 2023; Dong et al., 2023; Gulcehre et al., 2023; Yuan et al., 2024), with the general idea being to improve models with self-generated samples in an unsupervised or semi-supervised manner. He et al. (2019) is one early work in applying self-training to generative language models and points out the importance of introducing noises during pseudo-label generation to increase the sample diversity. In the large language model era, Gulcehre et al. (2023) propose Reinforced Self-Training (ReST), where they use a scoring function to select self-generated samples and augment the training data. Similarly, Yuan et al. (2024) proposes self-rewarding that scores samples with the LLM itself and trains the model with direct preference optimization (DPO) (Rafailov et al., 2024) on the scored samples. Self-training has also been employed to improve the chain-of-thought reasoning (Nye et al., 2022; Wei et al., 2022) ability of LLMs (Uesato et al., 2022). For example, Zelikman et al. (2022) propose to ask an LLM to generate rationales given questions and improve the LLM with its own generated reasoning. Re-ReST falls under the self-training paradigm, and different from previous work, our aim is to generate useful samples efficiently for self-training.
@@ 5 Conclusion
In conclusion, our study presents Reflection-Reinforced Self-Training (Re-ReST), an approach that improves self-training for language agents. Specifically, we integrate a reflection model to efficiently obtain high-quality samples for augmenting the self-training data. Our experiments demonstrate that Re-ReST outperforms self-training methods across various tasks, confirming the efficiency and effectiveness of incorporating a reflection mechanism.
Within the proposed framework, in the future, we can 1) improve the reflection mechanism, such as letting the reflector first learn principles from failures and then perform reflection efficiently (Zhang et al., 2024); 2) develop better training paradigms for the agent and reflector, such as jointly training the two models and letting them learn from each other instead of training them independently. We can also investigate its applications in other tasks and domains.

### v3

@@ 3.2 Sequential Decision-Making
We also assess the proposed approach on sequential decision-making using ALFWorld (Shridhar et al., 2021). ALFWorld comprises a collection of text-based settings designed to test an agent’s ability to complete multi-step tasks across diverse interactive environments. Following Yao et al. (2022); Shinn et al. (2023), we operate under the assumption that the agents are devoid of any access to successful trajectories, relying solely on a binary indicator of task success or failure. Our evaluation encompasses testing the agent across 134 previously unseen environments, spanning six diverse tasks. These tasks range from locating concealed items and transporting objects to interacting with objects using other items.
| Model | Sample Acc. | Success Rate |
| Few-Shot | - | 8.9 |
| Self-Training | 11.2 | 37.3 |
| Re-ReST | 48.0 | 51.4 |
@@
We build the agent and the reflector upon the Llama2-7b (Touvron et al., 2023). At each step, the agent can either contemplate its next move or generate admissible actions for execution as in Yao et al. (2022). Following the heuristics outlined by Shinn et al. (2023), we trigger the reflector model for self-reflection if the agent repeats an action with the same response over three cycles, or if it performs over 30 actions in an environment.
@@
We use one-shot prompting instead of the two-shot prompting in Shinn et al. (2023) for the models so that we can better fit a trajectory into the context window of Llama-2. We train the agent and reflector models on the collected trajectories for 2 epochs with a learning rate of 2e-5 using LoRA.
@@ Results.
As shown in Table 2, it is evident that the base Llama model faces challenges in adapting to the experimental environment, but self-training can significantly improve the model performance. A significant point to highlight is that the model operates without access to complete trajectories during the experiment. Despite this limitation, it demonstrates a notable improvement in performance within unseen environments—increasing the success rate from 8.9% to 37.3% through the utilization of self-augmented trajectories. Furthermore, the implementation of the reflector contributes a 14.1% uplift in success rates, which affirms the efficacy of our proposed method.
@@ 3.3 Programming: Code Generation and Visual Question Answering
For code generation, we experiment with the Python code writing task on MBPP (Austin et al., 2021) and visual programming on GQA (Hudson and Manning, 2019). The MBPP benchmark consists of around 1,000 Python programming problems, with each problem paired with unit test cases. We follow its official split for the training and test data. The availability of the training set and its provided unit test cases make it suitable for our reflector to reflect and correct the model-generated code. For GQA, we randomly sample a subset of 5,000 data points for training and 1,000 data for testing.
@@
We build both the agent model and the reflector upon the CodeLlama-13B model (Roziere et al., 2023). For MBPP, following Roziere et al. (2023), the agent model is given the unit test cases during code generation. Similarly, the reflection model is given the agent generation and its unit test results as the environmental feedback, and then generates a corrected version. For GQA, following Surís et al. (2023), we build the agent by providing a pre-defined set of visual APIs (e.g. object detection) and prompt the model to generate code using the APIs.
@@
For MBPP, we use zero-shot and three-shot prompting for zero-shot agent and reflector data generation. For GQA, we follow the prompt in Surís et al. (2023) for the model for sample generation. For both datasets, the agent model samples 3 generations per training instance as before. We do not use the provided ground truths for MBPP training for consistency with the other experimental settings. The agent and reflector models are trained for 3 epochs with a learning rate of 3e-4 using LoRA.
@@
| Sample Acc. | P@1 | Sample Acc. | Score | |
| Zero-Shot | - | 48.6 | - | 40.9 |
| Self-Training | 66.9 | 54.5 | 44.7 | 41.9 |
| Re-ReST | 77.3 | 56.4 | 55.7 | 42.6 |
@@ Results.
As in Table 3, for MBPP, because CodeLlama is trained on a large amount of code generation corpus, the base CodeLlama model can achieve a decent performance without any finetuning. The high pass rate results in many of the training instances being used for self-training. After self-training on the MBPP training data, the model performance can be improved from 48.6% to 54.5%. The reflector model can generate more self-training data and the pass rate can be improved with the reflector-generated data. For GQA, similar improvements can be seen, indicating that our method is also applicable in visual programming.
@@ 3.4 Text-to-Image Generation
| Model | Sample Acc. | VPEval Skill Score | ||
@@
| Model | Sample Acc. | EM |
@@
| S.C. (6 agents) | 30.8 |
| S.C. (3 agents + 3 reflectors) | 32.0 |
| Oracle (3 agents + 3 reflectors) | 36.8 |
| Model | HotpotQA EM | MBPP Pass@1 | VPEval Score | ||
@@
We use VPGen in Cho et al. (2023) as our base model, which is based on Vicuna-13B (Chiang et al., 2023) and is finetuned for text-to-layout generation on multiple constructed image-text datasets. The generated layouts are fed into an external model (i.e., GLIGEN (Li et al., 2023b)) for image generation. We build both the agent and reflector upon the VPGen model.
@@
We use VPGen to perform inference on their training data, and evaluate the generations using VPEval (Cho et al., 2023). Specifically, during evaluation, a visual question answering model (BLIP-2 (Li et al., 2023a)) is used to determine if the generated images correctly capture the input text information. The BLIP-2 generated results are treated as the environmental feedback for the reflector. We do not use zero-shot reflection results to train the reflector because LLMs cannot perform this task without finetuning. The agent and reflector are trained for 2 epochs with a learning rate of 1e-5 using LoRA.
@@ 3.5 Analysis
We investigate if we can simply sample more generations from the language agent for self-training and achieve comparable performance with our reflector-augmented method. Specifically, we try to sample generations for each instance, where is set to , and use the generated samples for self-training. As shown in Figure 3, if we keep sampling more generations from the language agent, the agent can indeed solve more instances and we can obtain an increasing amount of data for self-training. However, 1) the number of solved instances is still lower than the number of reflector-solved instances, demonstrating that the reflector can find the correct solutions more efficiently than sampling; 2) the model performance is not always improved with more training data and it cannot outperform our method even when trained with more generated samples, indicating that the quality of the self-training data is also important and our reflector can generate training data effectively for the agent.
@@
As illustrated, we propose to first train the reflector before using it to generate the self-training data. In this part, we investigate if we can use the reflector to perform self-correction in a zero-shot manner and then train the language agent. As in Table 5, we find that while the reflector can perform self-correction without any finetuning and improve the performance of the language agent, further improvements can be made if we specifically train the model for self-correction, demonstrating the effectiveness of our proposed reflector training strategy.
Test-Time Reflection without Ground-Truth Feedback.
Previously, our reflector functions only during training and is not used during inference because it is often impossible to obtain ground-truth feedback, which is required for reflection methods to work Huang et al. (2024). In this section, we propose employing self-consistency (Wang et al., 2023a) to enable test-time reflection and address this limitation. Self-consistency is a decoding technique that combines multiple model predictions by sampling various reasoning paths and then selecting the most consistent answer through a majority vote. This approach allows us to apply the reflector during inference. Specifically, we sample multiple answers from our model and perform reflection on each output, regardless of correctness. We then aggregate all the answers using self-consistency. As in Table 6, integrating our reflector with self-consistency (3 agent samples and 3 reflection samples) achieves improvements over baseline (self-consistency with 6 model samples). This demonstrates the potential application of our method during inference, overcoming the current limitation of requiring ground-truth feedback for reflection methods.
@@
Our reflector turns incorrect samples into correct ones, naturally making negative-positive pairs suitable for preference optimization objectives such as DPO. In this part, we investigate the application of DPO in our method. As in Table 7, integrating DPO into our method can generally improve or achieve comparable performance with training models only with supervised training on positive samples, indicating our compatibility with DPO.
@@ 4 Related Work
In this section, we first overview the research progress in language agents, then briefly describe self-training and self-correction methods for improving language agents. We also summarize the major differences between our work and previous language agent methods in Table 8.
@@
Language agents refer to language models that interact with the world in general. It has been demonstrated that LLMs can perform actions by generating specific commands (Nakano et al., 2021; Huang et al., 2022; Ahn et al., 2022) and calling external tool APIs (Lu et al., 2023; Schick et al., 2023; Gou et al., 2024). By integrating the model reasoning and acting abilities, ReAct (Yao et al., 2022) asks an LLM to first generate reasoning traces and then act accordingly, which is then improved by follow-up works through inference-time techniques such as reflection (Shinn et al., 2023) and planning (Yao et al., 2023; Yang et al., 2023). Recently, finetuning agents (Chen et al., 2023; Yin et al., 2024) have attracted attention from the research community. However, most of the existing works attempt to distill knowledge from a relatively strong LLM (e.g., GPT-4) to a weaker LLM (e.g., LLaMa-2). By contrast, our work bootstraps a language agent’s performance by utilizing its own reflective ability without using external models.
@@
Various self-training algorithms have been proposed to improve language models (He et al., 2019; Huang et al., 2023; Dong et al., 2023; Gulcehre et al., 2023; Yuan et al., 2024), with the general idea being to improve models with self-generated samples in an unsupervised or semi-supervised manner. He et al. (2019) is one early work in applying self-training to generative language models and points out the importance of introducing noises during pseudo-label generation to increase the sample diversity. In the large language model era, Gulcehre et al. (2023) propose Reinforced Self-Training (ReST), where they use a scoring function to select self-generated samples and augment the training data. Similarly, Yuan et al. (2024) proposes self-rewarding that scores samples with the LLM itself and trains the model with direct preference optimization (DPO) (Rafailov et al., 2023) on the scored samples. Self-training has also been employed to improve the chain-of-thought reasoning (Nye et al., 2022; Wei et al., 2022) ability of LLMs (Uesato et al., 2022). For example, Zelikman et al. (2022) propose to ask an LLM to generate rationales given questions and improve the LLM with its own generated reasoning. Re-ReST falls under the self-training paradigm, and different from previous work, our aim is to generate useful samples efficiently for self-training.
@@ 5 Conclusion
Our study studies the applications of self-training in language agents and improves it with Reflection-Reinforced Self-Training (Re-ReST), an approach that efficiently obtains high-quality samples for self-training with a reflector. Our experiments demonstrate that Re-ReST outperforms self-training methods across various tasks, confirming the efficiency and effectiveness of incorporating a reflection mechanism. Within the proposed framework, in the future, we can improve the reflection mechanism and develop better training paradigms for the agent and reflector.
Acknowledgement
We thank anonymous reviewers and UCLA NLP group members for their insightful feedback. This research is based upon work supported by DARPA ECOLE Program No. #HR00112390060, OFFICE OF NAVAL RESEARCH Award #N00014-23-1-2780, an Amazon AGI foundation research award, a google research scholar grant, and CISCO sponsored research award.
Limitations
Our approach is predicated on the availability of ground-truth feedback during the training process. While this assumption holds true for many language agent tasks, it presents challenges when applied to broader contexts. Specifically, acquiring accurate ground-truth feedback can be difficult in diverse, real-world scenarios. This limitation underscores a key aspect of our study: it is primarily concentrated on language agent tasks, thereby neglecting the potential applications and implications within the broader scope of general language modeling. This suggests the need for future research to explore and address the complexities of applying our methods to general language modeling tasks, where ground-truth feedback may not be as readily accessible or reliable. Another potential risk of the method is that through self-training, the biases encoded in LLMs can be amplified, and careful calibrations should be conducted before the deployment of our method.